OpenStudy (anonymous):

A mass m = 11.0 kg rests on a frictionless table and accelerated by a spring with spring constant k = 5008.0 N/m. The floor is frictionless except for a rough patch. For this rough path, the coefficient of friction is μk = 0.46. The mass leaves the spring at a speed v = 2.8 m/s. 3) The mass is measured to leave the rough spot with a final speed vf = 1.5 m/s. How much work is done by friction as the mass crosses the rough spot? What is the length of the rough spot? In a new scenario, the block only makes it (exactly) half-way through the rough spot. How far was the spring compressed fro

OpenStudy (anonymous):

sorry, saw you were off line so thought you were gone for the evening...

OpenStudy (anonymous):

let me look it over..

OpenStudy (anonymous):

sure, so... work done by the rough spot (friction) = 1/2*11*( 1.5^2 -2.8^2)

OpenStudy (anonymous):

length of the rough spot is from Work = F*d = mu*mg*d

OpenStudy (anonymous):

now the last part should be clearer... it's a little tricky but not too tough...

OpenStudy (anonymous):

find W= mu*mg*d/2 that's the KE the mass had when it encountered the patch (and also the KE when it left the spring) KE=PEspring mu*mg*d/2 = 1/2 *k*x^2

OpenStudy (anonymous):

thank you soo much ur a life savor litterally..i have 2 more questions i cant get .if you wouldnt mind helping me on those id appreciate it soo much

OpenStudy (anonymous):

work energy theorem problems? you should be able to do them now... I'll answer any specific questions you have about particular details...

OpenStudy (anonymous):

now i have conservative forces and potential energy problem